The width of a rectangle is 5 less than twice its length. If the area of the rectangle is 23 cm^2, what is the length of the diagonal?

Answer by stanbon(75887)

The width of a rectangle is 5 less than twice its length. If the area of the rectangle is 23 cm^2, what is the length of the diagonal?
---
Length:: x
Width:: 2x-5
-----
x(2x-5) = 23
2x^2 -5x - 23 = 0
x = 4.86
2x-5 = 4.72
-----
Ans: diag = sqrt(4.86^2+4.72^2) = 6.77 cm
------------
Cheers,
Stan H.
-------------